You are offered an investment opportunity with the guarantee that your investment will double in 5 years. Assuming annual compounding, what annual rate of return would this investment provide?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value($2x)
P=present value($x say)
r=rate of interest
n=time period.

2x=x*(1+r/100)^5

(2)^(1/5)=(1+r/100)

(1+r/100)=1.1487

r=(1.1487-1)*100

=14.87%(Approx).